OpenStudy (cutiecomittee123):

Find the exact value of each o the following expressions using trigonometric sum identities cos75 degrees sin195 degrees tan105 degrees

OpenStudy (flvskidd):

cos .2588 sin -.2588 tan -3.7

OpenStudy (cutiecomittee123):

how?

OpenStudy (flvskidd):

That's just my calculator I don't know what they mean by trigonometric sum identities.

OpenStudy (cutiecomittee123):

Thanks then:)

OpenStudy (flvskidd):

here i'll look it up

OpenStudy (flvskidd):

this might help you https://www.youtube.com/watch?v=ImUluLTJFno

OpenStudy (cutiecomittee123):

yeah it helped with the first one

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you are supposed to use the addition angle formula for this one

OpenStudy (misty1212):

\[75^\circ=30^\circ+45^\circ\] and \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]

OpenStudy (misty1212):

put in \(a=30^\circ\)and \(b=45^\circ\) compute the result

OpenStudy (cutiecomittee123):

Yes I have solved that one and I kinda understand but once it gets to using the unit circle I am a little lost

OpenStudy (misty1212):

you need to know how to find \(\cos(30)\) and so on?

OpenStudy (cutiecomittee123):

On the unit circle, yes

OpenStudy (misty1212):

find the point corresponding to the angle of 30 on the unit circle let me know when you are looking at \((\frac{\sqrt3}{2}, \frac{1}{2})\)

OpenStudy (cutiecomittee123):

yes, okay that helps me a lot!!! thanks for that

OpenStudy (misty1212):

the first coordinate is cosine , the second is sine

OpenStudy (misty1212):

so \[\cos(30)=\frac{\sqrt3}{2}\]and \[\sin(30)=\frac{1}{2}\]

OpenStudy (cutiecomittee123):

okay so I see that I just plug those in and then simplify.

OpenStudy (misty1212):

yes

OpenStudy (cutiecomittee123):

For the second one I dont know if I should use sin(a-b) or sin(a+b)

OpenStudy (cutiecomittee123):

and I know its 195, so should I do 180+15? Or should I use 210-15?

OpenStudy (cutiecomittee123):

@misty1212

OpenStudy (cutiecomittee123):

@jim_thompson5910 can you help me?